OpenStudy (anonymous):

find the 10th term from the end of the ap 4,9,14.....254

OpenStudy (aravindg):

what does you see as the similarity between consecutive terms?

OpenStudy (aravindg):

4,9,14.....what the relation between each consecutive term ?

OpenStudy (aravindg):

@aaliya reply?

OpenStudy (anonymous):

2 numbers

OpenStudy (aravindg):

whats the relation between the 2 numbers?

OpenStudy (anonymous):

i do not understand

OpenStudy (aravindg):

4,9,14 ...what kind of progression is this ?

OpenStudy (aravindg):

Is it AP ,GP or HP ?

OpenStudy (anonymous):

ap

OpenStudy (aravindg):

well done :)

OpenStudy (aravindg):

now you need 10th term from end

OpenStudy (anonymous):

tell me answer i have less time..pls help me

OpenStudy (aravindg):

ok then find what will be n value for 254

OpenStudy (aravindg):

a+(n-1)d=254

OpenStudy (aravindg):

once you get that subtract 10 from it

Parth (parthkohli):

Just another way to think about it: 9 is the second term. 19 is the fourth term. 29 is the sixth term. . . .

OpenStudy (aravindg):

can you do that much ?

OpenStudy (anonymous):

ok u give me full solution i understand it .then i have promble i ask u

OpenStudy (aravindg):

Sorry Openstudy doesnt work like that ..we guide you through the answering process rather than giving full solution. .Refer Openstudy COC

OpenStudy (aravindg):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

please fast i have 34 question i have only 2.00 clock time pls........

Parth (parthkohli):

i = -1; while (i < 254) { i += 5; print i }

OpenStudy (aravindg):

@aaliya ya u can do it in just 1 minute ....follow what i have said above

Parth (parthkohli):

I can do it in 2 seconds by the method I posted: 9 is the second term. 19 is the fourth term. 29 is the sixth term . . .

OpenStudy (anonymous):

ok parth

Parth (parthkohli):

\[\textbf{HINT}\]The answer is the solution to this integral:\[-10+ \frac{64}{\pi^3} \int_0^\infty \frac{ (\ln x)^2 (15-2x)}{(x^4+1)(x^2+1)}\ dx\]

OpenStudy (aravindg):

wow that requires only fraction of a second !

Parth (parthkohli):

Indeed, you just have to look at the integral.

Parth (parthkohli):

@aaliya Connections Academy?

OpenStudy (aravindg):

All the best @aaliya Since you are not willing to take my help . :)

OpenStudy (anonymous):

sorry.........